# Additive manufacturing of a 3D-segmented plastic scintillator detector for tracking and calorimetry of elementary particles

### v1

@@ Abstract
Plastic-scintillator detectors are among the most common devices used for the detection of elementary particles. They provide good particle identification combined with excellent time resolution, whilst being inexpensive due to the affordability of plastic materials. Particle tracking is achieved by segmenting the scintillator into smaller, independent, optically-isolated voxels. Enhancing the performance of future particle detectors necessitates larger total volumes, possibly combined with even finer segmentation. However, manufacturing such designs with current production strategies is challenging. These strategies involve a variety of time-consuming and costly fabrication processes, followed by the assembly of millions of individual parts. The difficulty in scaling up such a complex workflow underscores the need for technological advancements, which can be met by additive manufacturing. This method enables the construction of complex, monolithic geometries in a single operation. These geometries consist of fine three-dimensional granular sub-structures and require the integration of multiple types of plastic materials, as well as space to accommodate optical fibers, all within a compact volume of several cubic meters.
This article presents the fabrication and performance evaluation of the first additive manufactured single-block plastic scintillator detector. To achieve this, a new method called Fused Injection Modeling has been specially developed. The detector is capable of three-dimensional tracking of elementary particles and accurately measuring their stopping power. Its performance is comparable to the current state of the art of plastic scintillator detectors. This work paves the way towards a new feasible, time and cost-effective process for the production of future scintillator detectors, regardless their size and difficulty in geometry.
@@ I Introduction
Plastic scintillator (PS) detectors, invented in the early 1950s [1], are widely used in the detection of elementary particles in high-energy physics, nuclear physics, astroparticle physics, as well as in many applications like muon tomography, proton computed tomography for hadron therapy, fast-neutron detection and non-destructive imaging.
@@
A PS is an organic material composed of a mixture of carbon and hydrogen-based molecules, typically polystyrene (PST) or polyvinyltoluene (PVT). Molecules of an activator like p-terphenyl (pTP), 2,2-p-phenilene-bis(5-pheniloxazole) (POPOP), 2,5- diphenyloxazole (PPO) are introduced into the polymer at a level between a few per mille or percent by weight. Such formula has not undergone any major change in the last fifty years. The scintillation mechanism consists of a few steps, highlighted in Fig. 1. When a charged particle propagates through a PS, the molecules of the polymer matrix get excited. A fast short-range resonant non-radiative dipole-dipole interaction, called Förster mechanism [9], efficiently transfers the excitation energy to the activator, which de-excites and emits near-ultraviolet (UV) photons with a minimal emission delay. A second dopant, called a shifter, is usually added to change the wavelength of the light to avoid absorption in the material. More details about the scintillation mechanism in organic materials can be found in [10, 11]. Often, the light produced in PS is collected by wavelength-shifting (WLS) fibers. These fibers shift the light from the violet/blue spectrum, which is the typical emission range of PSs, to green, where the attenuation length is maximum. They then guide it towards photodetectors, benefiting from the over several meters low attenuation length of the fiber [4].
Figure 1: Top: representation of the scintillation process in PS doped with a primary and a secondary fluor when the material is crossed by a charged particle. Bottom left: emission spectrum of a 3D printed PS sample (blue) taken with a spectrophotometer, Kuraray Y11 WLS fiber absorption (green) and emission (red) spectra. Bottom right: measured transmittance of white reflector sheets with the same thickness but different print orientation compared to the printing bed. Vertical (side) walls in red and horizontal (top and bottom) walls in blue of a PS voxel.
@@
PS detectors with tracking capabilities are made of scintillating voxels, which are optically isolated from each other. The typical light reflector is composed of a white diffuser, such as paint containing titanium dioxide ($\text{TiO}_{2}$), white polytetrafluoroethylene (PTFE), or is created using a process of chemical etching [17]. Co-extrusion of PS and optical reflector has also been conducted [6, 4]. Another option with good performance consists of wrapping each scintillating element with Tyvek [18]. However, such a solution is unfeasible for detectors with a large number of PS elements [19, 20, 21].
[…]
@@
Attempts to simplify detector fabrication were reported in Ref. [30], where a prototype of PS cubes glued together was obtained with a tolerance of about 200 $\mu$m. However, such a method is not feasible for the production of a single 3D volume of PS cubes, but only of 2D layers, whose production would also be time consuming for large-scale applications. The above considerations call for the development of a novel manufacturing process that allows for the easy production of several thousand optically-isolated PS cubes in a single block of plastic. Our solution takes advantage of additive manufacturing (AM), which opens the door to new automated processes that could drastically simplify the construction of future particle detectors. More commonly known as rapid prototyping or 3D printing, AM processes build the designed parts through a layer-by-layer addition of new material. AM technologies are capable of fabricating customizable, monolithic parts with multiple materials and complex internal geometries, and can greatly reduce production time and cost. Two common AM technologies for polymers are: stereolithography (SLA) [31], where liquid resin is solidified after curing with, for example, ultraviolet (UV) light; and Fused Deposition Modeling (FDM) [32], where, analogous to extrusion, a thermoplastic material in the form of a thin wire is passed through a feeding into a melting system and deposited on a print bed line-by-line with the help of a nozzle tip [33, 34]. 3D printing of PS with SLA has been reported in literature. However, studies have shown the necessity of either developing a new chemical composition or binding PS granules into a polymer matrix. This poses a challenge in achieving competitive performance levels in terms of acceptable light yield and attenuation length compared to standard PS [35]. Furthermore, it is not an easy task to achieve 3D printing of multiple materials, along with the production of hollow objects that possess smooth and consistent inner surfaces, which is needed for the insertion of WLS fibers.
As we have shown in Ref. [36], the most promising additive manufacturing option is FDM. It allows the use of the same chemical composition as standard scintillators in the form of a filament, ensures high transparency and multi-material printing. For more details on the FDM process and the naming convention of its components, we refer to Sec. IV.2. Furthermore, we have demonstrated that FDM can successfully 3D print PST-based scintillators. These scintillators are extensively utilized in particle detectors and exhibit a light yield performance comparable to the current state of the art. Another independent work reached a similar conclusion [37]. Later, we showed a technical attenuation length of about 19 cm measured in a bar of 5 cm in length, sufficient for fine-granularity scintillator detectors [38].
@@
The absorption and emission spectra of the 3D printed PS are shown in Fig. 1. For such a scintillator composition, the emission spectrum of the 3D printed PS peaks around 420 nm upon the excitation with 386 nm light. This result is consistent with the UPS-923A standard scintillator reported in [39]. It also shows that the measured scintillation spectrum closely matches the absorption spectrum of green WLS fibers such as the Kuraray Y11 [40], allowing for an efficient readout of the scintillation light. In Ref. [38], the first 3D printed matrix of PS cubes optically isolated with a custom-fabricated white reflective filament (PST, or polymethyl methacrylate (PMMA)) mixed with $\text{TiO}_{2}$ was produced and showed a low cube-to-cube light crosstalk, less than 2%. However, after 3D printing, no work reported in literature could avoid the use of subtractive processes, like polishing the outer surface to achieve a good geometrical tolerance. Moreover, none of the studies ever attempted to 3D print hollow PS objects to host WLS fibers.
@@ II Results
A Computational Fluid Dynamics (CFD) analysis was performed to determine the material requirements of the melting components, heat block and nozzle; the heat shielding parts, feeding tube and heat break; combined with the process parameters, heat block temperature, extrusion speed, and polymer temperature at the nozzle orifice. A detailed explanation of the nomenclature of the machine components, design and mode of operation of the manufacturing process is found in Sec. IV.2.
@@
The fabrication procedure, using the custom manufactured modified extrusion system (Fig. 2, left, bottom), contained three main steps and is described in more detail in Sec. IV: the fabrication of one matrix-layer of the reflective frame via FDM; the insertion of metal rods into the voxel cavity through already 3D-printed holes in the frame; followed by the forming of the cube-shaped volume with melted scintillation material (Fig. 2, middle). These processes can be repeated until a SuperCube of the desired size has been obtained. Finally, the WLS fibers were placed through the PS voxel via the cylindrical holes produced by the metal rods (Fig. 2, top right).
The optically reflective frame constituted the mold with which the PS was formed and thus needed to withstand the pressure and heat from the extruded melt pool. As reported in our previous work [38], in order to manufacture a 3D printed layer of optically-isolated cubes, a custom optical reflector filament was fabricated with either PMMA or PST. A transmittance of less than 10% and an optical cube-to-cube light crosstalk of less than 2% were achieved, the material could not maintain its structural integrity during the injection of the PS. This was due to the similar heat resistance of the optical reflector compared to the PS, with the consequence of swelling and wall bending during filling. Hence, for the FIM method, a white, more heat-resistant polycarbonate-PTFE filament was used, which combines good optical properties of PTFE and a heat-resistance significantly higher than the previously used materials.
In FDM, vertical walls are created by stacking lines on top of each other, while horizontal walls are formed by placing adjacent lines on the same print plane. This results in different fill factors, thus different light transmissive properties. The corresponding transmittance for a wavelength of 420 nm, measured with a monochromatic light source, resulted in 13% for horizontally and 18% for vertically built walls, as shown in Fig. 1. Consequently, to obtain a uniform cube-to-cube light crosstalk, horizontal walls were fabricated with a thickness of 1.2 mm and the more transmissive vertical walls with 1.5 mm. The more heat-resistant reflective material retained its as-built cube shape throughout the filling fabrication of the PS, as can be seen in Fig. 2, bottom right. Caliper measurements of vertical walls showed an average thickness of $t_{\text{walls, mean}}$ = 1.51 mm, deviating by only 0.01 mm from the nominal $t_{\text{walls, designed}}$ = 1.5 mm with a standard deviation of $t_{\text{wall, std}}$ = 0.01 mm. Both the retention of the reflective frame geometry after filling and the accurate fabrication of the wall thickness ensured a consistent active volume of scintillation material in every voxel. More details about the fabrication of the white reflector walls can be found in Sec. IV.2.1.
@@
Figure 3: Left: SuperCube illuminated with UV light without white reflector on its top face to show the inner structure of the optically-isolated PS cubes crossed by WLS fibers. Right: Completed SuperCube manufactured with FIM. All six faces are covered with white reflector. No post-processing was necessary.
@@
The response of the SuperCube was characterised with cosmic particles in terms of single-cube scintillation light yield and cube-to-cube scintillation light crosstalk, which are the main parameters that determine the detector performance in terms of particle tracking, identification and calorimetry. Cosmic particles entering the prototype from the top side and exiting through the bottom were selected for the analysis. Being mostly minimum ionizing particles (MIP), their typical energy deposited in PS is approximately 1.8 MeV/cm, which allows to precisely estimate the number of photons detected per unit energy loss. Moreover, with an energy spectrum that spans from a few hundred MeV to several GeV, cosmic particles typically produce through-going tracks in the detector with a very uniform energy deposition. Two selected cosmic particle events are shown in Fig. 4. By measuring the number of PE in all the readout channels and geometrically matching the two detector readout, the 3D track of the particle can be easily reconstructed and the number of scintillation photons, produced in each cube, measured. As shown in Fig. 5, the scintillation light yield of the 3D printed prototype was found to be comparable to that obtained with the one produced by cast polymerization (using the same scintillator composition) in [30]. A most probable value (MPV) of about 29 PE was measured for the 3D printed SuperCube, close to the MPV of the prototype produced with cast polymerization (about 28 PE).
Figure 4: Top: FIM manufactured SuperCube prototype instrumented with WLS fibers and SiPMs. Bottom: events of two detected cosmic particles crossing the SuperCube from top to bottom with a detected diagonal (left) and vertical (right) track. The 2D projections show the number of photons detected in each readout channel (color scale), while the 3D voxels show the reconstructed 3D particle track. In the matching of the 2D projections channels with less then 7 PE, likely from cube-to-cube light crosstalk, were not considered. The red line indicates a fitted particle trajectory.
@@ III DISCUSSION
The FDM manufactured PS showed a sufficient transparency for centimeter-sized voxels, but the FIM fabricated samples visibly improved the quality by reducing imperfections created by the merger of different deposited lines during the FDM process and the reduction of air bubbles within the PS volume. The improved transparency results from the straight down-to-up filling pattern, which creates a single melt pool that expands and forms the PS structure within the optical reflective frame. Multiple converging melt pools introduce non-uniformities that reduce the transparency of the PS. Additionally, the forming of a cubic-centimeter sized PS cube with FIM needs around 30 seconds, whereas the same structure takes around 5 minutes using FDM fabrication, i.e., about ten times as long. The faster per-unit manufacturing time is crucial when it comes to the production of large volume, highly-segmented particle detectors that can include millions of isolated voxels.
@@
The geometrical tolerance of the optically isolating frame produced by FIM is improved compared to the particle detector manufactured solely by FDM. This is a result of the introduction of a heat-resistant material that is able to preserve its designed shape during the filling of melted PS, thus creates a distinct boundary between the individual PS voxels. In FDM, both the scintillation and the reflective filament are made of plastics with similar thermal characteristics, which can result in a mixing at the boundary layers between these two materials and also a warping and distortion of the reflective geometry. Although the tolerance cannot be considered as good as the one achievable with injection molding or cast polymerization (better than $50~{}\mu\text{m}$ for plastics), it is worth noting, that it is affected by the type of 3D printed material, the printing strategy and the precision of the machine, which will be improved in future studies.
@@
The largest advantage of the FIM method compared to FDM or more traditional methods is the capability of creating precise, close-fitting holes through the entire particle detector to host WLS fibers. FDM is notorious for producing poor quality small holes in terms of size consistency, shape and surface roughness, which results in a low WLS fiber to PS contact area, especially perpendicularly to the print plane. Injection molding has the capability to create hollow components within the manufactured structure. However, incorporating a white reflector without causing any harm to the pre-existing holes poses a significant challenge. In our next prototypes also a third vertical hole will be added.
@@
In terms of the detection performance, the quantity of scintillation light, collected from the FIM manufactured PS cubes crossed by cosmic particles, is similar to that observed from PS cubes produced with traditional methods like cast polymerization. The cube-to-cube crosstalk at the few-percent level is comparable to that of typical PS detectors and is low enough to provide unambiguous 3D particle tracking.
@@
The light crosstalk is about four times higher than with the cast polymerization prototype, due to the higher wall transmittance. Hence, one would expect an approximately 20% lower total light yield. However, this was not observed, as shown in Fig. 5. One may conclude that the additive manufactured PS has an intrinsic light output higher than the one from cast polymerization, but this was not the result of our past work [38, 36]. It is well known, that the light yield of a PS detector can be enhanced by improving the light trapping efficiency of the WLS fibers through the reduction of air gaps, that is the increase of the contact area between the PS and the WLS fibers, whose refractive indices are very similar and thus allows a high transmittance from the scintillation material to the fiber. This can be obtained by using optical grease [41, 42, 43], which is difficult to homogeneously introduce in long, small-diameter holes. Using the FIM method, it was possible to create holes that span the length of the SuperCube of only 100 $\mu$m larger than the WLS fibers, resulting in a large contact surface between the PS cubes and the fibers. This feature is expected to increase the total light yield and compensate for the light loss from the observed crosstalk.
@@
In conclusion, we have successfully achieved the additive manufacturing of a 3D optically-segmented plastic scintillator detector. To the best of our knowledge, this is the first time such a feat has been accomplished. This detector is capable of both particle tracking and energy loss measurement. It demonstrates a tool that opens new possible concepts of particle detectors characterised by a very fine 3D granularity in very large active volumes, above the tonne-scale. Although not tested in this work, millimeter-sized granularity and bigger units, e.g. $30\times 30\times 30~{}\text{cm}^{3}$ or above, are expected to be achieved without problems. We believe that this new technology opens up new possibilities for the future of particle physics experiments. In particular, long-baseline accelerator neutrino oscillations [44, 45, 46, 47] or those searching for new neutrino sterile states via short-baseline oscillations [48, 49, 50, 51, 52, 53] can profit of the developed additive manufacturing process by building very large but finely-segmented scintillator detectors to obtain a large sample of very detailed images of neutrino interactions. Moreover, the developed production method will allow to easily build highly-performing fine-granularity electromagnetic or hadronic sampling calorimeters, enabling high-resolution particle flow analysis and fulfill the requirements of future collider experiments [22, 54]. FIM can be performed with various types of filaments, where the white reflector could be covered by an additional layer made of heavier nuclei or doped material. Finally, large-volume 3D detector geometries, made possible by the FIM process, can allow for a high-efficiency detection of fast neutrons with a precise measurement of their time of flight, thanks to the large content of hydrogen in PS and its short decay time [19, 26]. This is of fundamental importance for next generation neutrino physics experiments, providing constraints on leading systematic uncertainties [25, 55, 24, 56].

### v2

@@ Abstract
Plastic-scintillator detectors are devices used for the detection of elementary particles. They provide good particle identification with excellent time resolution, whilst being inexpensive due to the affordability of plastic materials. Particle tracking is achieved by segmenting the scintillator into smaller optically-isolated 3D granular sub-structures which require the integration of multiple types of plastic materials as well as several thousands of tiny holes through a compact volume of several cubic meters. Future particle detectors necessitate larger volumes, possibly with even finer segmentation. However, manufacturing such geometries with current production strategies is challenging, as they involve time-consuming and costly fabrication processes, followed by the assembly of millions of individual parts. The difficulty in scaling up such a workflow can be addressed by additive manufacturing, enabling the construction of complex, monolithic geometries in a single operation. This article presents the fabrication of the first additive manufactured plastic scintillator detector, capable of 3D tracking elementary particles and measuring their stopping power. Its performance is comparable to the state of the art of plastic scintillator detectors. This work paves the way towards a new feasible, time and cost-effective process for the production of future plastic-based scintillator detectors, regardless their size and difficulty in geometry.
@@ I Introduction
Plastic scintillator (PS) detectors, invented in the early 1950s [1], are widely used in the detection of elementary particles in high-energy physics (HEP) [2, 3, 4, 5, 6], nuclear physics [7], astroparticle physics [8, 9], as well as in many applications like muon tomography [10], proton computed tomography for hadron therapy [11], fast-neutron detection [12, 13] and non-destructive imaging [14].
@@
A PS is an organic material composed of a mixture of carbon and hydrogen-based molecules, typically polystyrene (PST) or polyvinyltoluene (PVT). Molecules of an activator like p-terphenyl (pTP), 2,2-p-phenylene-bis(5-phenyloxazole) (POPOP), 2,5-diphenyloxazole (PPO) are introduced into the polymer at a level between a few per mille and percent by weight. The scintillation mechanism consists of a few steps, highlighted in Fig. 1. When a charged particle propagates through a PS, the molecules of the polymer matrix get excited. A fast short-range resonant non-radiative dipole-dipole interaction, called Förster mechanism [18], efficiently transfers the excitation energy to the activator, which de-excites and emits near-ultraviolet (UV) photons with a minimal emission delay. A second dopant, called shifter, is usually added to change the wavelength of the light to avoid absorption in the material. More details about the scintillation mechanism in organic materials can be found in [19, 20]. In HEP experiments, the light produced in PS is frequently collected using wavelength-shifting (WLS) fibers. These fibers shift the light from the violet/blue spectrum, which is the typical emission range of PS, to green, where the attenuation length is longer. Then, they guide it towards photodetectors, taking advantage of the fiber’s long attenuation length, which spans over several meters [2].
Figure 1: Top: representation of the scintillation process in PS doped with a primary and a secondary fluor when the material is crossed by a charged particle. Bottom: emission spectrum of a 3D printed PS sample (blue) taken with a spectrophotometer, Kuraray Y11 WLS fiber absorption (green) and emission (red) spectra.
@@
PS detectors composed of scintillating voxels, optically isolated from each other, are optimal for particle tracking. The typical light reflector is composed of a white diffuser, such as paint containing titanium dioxide ($\text{TiO}_{2}$), white polytetrafluoroethylene (PTFE), or is created using a process of chemical etching [26]. Co-extrusion of PS and optical reflector has also been conducted [4, 2]. Another option with good performance consists of wrapping each scintillating element with Tyvek [27]. However, such a solution is unfeasible for detectors with a large number of PS elements [28, 29, 30].
In recent years, there have been advancements in the development of novel three-dimensional (3D) granular scintillating detectors for imaging electromagnetic and hadronic showers [31], as well as neutrino interactions [32, 28, 29, 30]. Combining these geometries with the PS sub-ns response is ideal for efficient neutron detection with kinetic energy reconstruction by time-of-flight [33, 34, 35, 36]. For instance, a neutrino detector made of two-million PS cubes, each of size 1 cm3, with a total active mass of two tonnes, has been proposed [28], prototyped [29, 30], built and started collecting data as of 2023 at the T2K neutrino experiment in Japan [37, 38]. Simulation studies conducted independently on a comparable 3D granularity plastic scintillator detector demonstrated that a tracking resolution ranging between 2 mm and 4 mm, depending on the utilized reconstruction algorithm, can be attained for minimum ionizing particles such as muons [39]. The complexity in geometry of such a detector requires multiple vastly different manufacturing steps including the fabrication of every single PS cube, optical isolation via chemical etching and drilling of the holes for WLS fiber placement. Additionally, the assembly of two-million cubes demands a significant amount of effort [40] and needs to be combined with a robust box that mechanically maintains the structural integrity of the entire construction [37].
@@
Attempts to simplify detector fabrication were reported in Ref. [41], where a prototype of PS cubes glued together was obtained with a tolerance of about 200 $\mu$m. However, such a method is not feasible for the production of a single 3D volume of PS cubes, but only of 2D layers, whose production would also be time consuming for large-scale applications. The above considerations call for the development of a novel manufacturing process that allows for the easy production of several thousand optically-isolated PS cubes in a single block of plastic. Our solution takes advantage of additive manufacturing (AM), which opens the door to new automated processes that could drastically simplify the construction of future particle detectors. More commonly known as rapid prototyping or 3D printing, AM processes build the designed parts through a layer-by-layer addition of new material. AM technologies are capable of fabricating customizable, monolithic parts with multiple materials and complex internal geometries, and can greatly reduce production time and cost. Two common AM technologies for polymers are: stereolithography (SLA) [42], where liquid resin is solidified after curing with, for example, ultraviolet (UV) light; and Fused Deposition Modeling (FDM) [43], where, analogous to extrusion, a thermoplastic material in the form of a thin wire is passed through a feeding into a melting system and deposited on a print bed line-by-line with the help of a nozzle tip [44, 45]. 3D printing of PS with SLA has been reported in literature. However, studies have shown the necessity of either developing a new chemical composition or binding PS granules into a polymer matrix. This poses a challenge in achieving competitive performance levels in terms of acceptable light yield and attenuation length compared to standard PS [46]. In recent years, significant progress has been made in the development of curable resins, primarily targeting applications beyond HEP. These advancements have demonstrated good performance in terms of light output, pulse shape discrimination, thermal neutron sensitivity, and various other properties when compared to commercially available alternatives [47, 48, 49, 50, 51, 52]. Anyhow, it is not an easy task to achieve 3D printing of multiple materials, along with the production of hollow objects that possess smooth and consistent inner surfaces, which is needed for the insertion of WLS fibers. As we have shown in Ref. [53], a promising additive manufacturing option is FDM. It allows the use of the same chemical composition as standard scintillators in the form of a filament, ensures high transparency and multi-material printing. For more details on the FDM process and the naming convention of its components, we refer to Sec. IV.2. Furthermore, we have demonstrated that FDM can successfully 3D print PST-based scintillators. Another independent work reached a similar conclusion [54]. Later, we showed a technical attenuation length of about 19 cm measured in a bar of 5 cm in length, sufficient for fine-granularity scintillator detectors [55].
@@
The absorption and emission spectra of the 3D printed PS are shown in Fig. 1. For such a scintillator composition, the emission spectrum of the 3D printed PS peaks around 420 nm upon the excitation with 386 nm light. This result is consistent with the UPS-923A standard scintillator reported in [56, 57]. It also shows that the measured scintillation spectrum closely matches the absorption spectrum of green WLS fibers such as the Kuraray Y11 [58], allowing for an efficient readout of the scintillation light. In Ref. [55], the first 3D printed matrix of PS cubes optically isolated with a custom-fabricated white reflective filament (PST, or polymethyl methacrylate (PMMA)) mixed with $\text{TiO}_{2}$ was produced and showed a low cube-to-cube light crosstalk, less than 2%. However, after 3D printing, no work reported in literature could avoid the use of subtractive processes, like polishing the outer surface to achieve a good geometrical tolerance. Moreover, none of the studies ever attempted to 3D print hollow PS objects to host WLS fibers.
@@ II Results
A Computational Fluid Dynamics (CFD) analysis was performed to determine the material requirements of the melting components, heat block and nozzle; the heat shielding parts, feeding tube and heat break; the process parameters, heat block temperature, extrusion speed, and polymer temperature at the nozzle orifice. A detailed explanation of the nomenclature of the machine components, design and mode of operation of the manufacturing process is found in Sec. IV.2.
@@
The fabrication procedure, using the custom manufactured modified extrusion system (Fig. 2, left, bottom), contained three main steps and is described in more detail in Sec. IV: the fabrication of one matrix-layer of the reflective frame via FDM; the insertion of metal rods into the voxel cavity through already 3D-printed holes in the frame; followed by the forming of the cube-shaped volume with melted scintillation material (Fig. 2, middle). These processes can be repeated until a SuperCube of the desired size has been obtained. Finally, the WLS fibers were placed through the PS voxel via the cylindrical holes produced by the removed metal rods (Fig. 2, top right). Metal rods were essential because WLS fibers could not be positioned during the PS forming process without risking thermal damage.
The optically reflective frame constituted the mold with which the PS was formed and thus needed to withstand the pressure and heat from the extruded melt pool. As reported in our previous work [55], in order to manufacture a 3D printed layer of optically-isolated cubes, a custom optical reflector filament was fabricated with either PMMA or PST. Although a transmittance of less than 10% and an optical cube-to-cube light crosstalk of less than 2% were achieved, the material could not maintain its structural integrity during the injection of the PS. This was due to the similar heat resistance of the optical reflector compared to the PS, with the consequence of swelling and wall bending during filling. Hence, for the FIM method, a white, more heat-resistant polycarbonate-PTFE filament was used, which combines good optical properties of PTFE and a heat-resistance significantly higher than the previously used materials.
In FDM, vertical walls are created by stacking lines on top of each other, while horizontal walls are formed by placing adjacent lines on the same print plane. This results in different fill factors, thus different light transmissive properties. The corresponding transmittance for a wavelength of 420 nm, measured with a monochromatic light source in air, resulted in 13% for horizontally and 18% for vertically built walls, as shown in Fig. 8. Consequently, to obtain a uniform cube-to-cube light crosstalk, horizontal walls were fabricated with a thickness of 1.2 mm and the more transmissive vertical walls with 1.5 mm. The more heat-resistant reflective material retained its as-built cube shape throughout the filling fabrication of the PS, as can be seen in Fig. 2, bottom right. Caliper measurements of vertical walls showed an average thickness of $t_{\text{walls, mean}}$ = 1.51 mm, deviating by only 0.01 mm from the nominal $t_{\text{walls, designed}}$ = 1.5 mm with a standard deviation of $t_{\text{wall, std}}$ = 0.01 mm. Both the retention of the reflective frame geometry after filling and the accurate fabrication of the wall thickness ensured a consistent active volume of scintillation material in every voxel. More details about the fabrication of the white reflector walls can be found in Sec. IV.2.1.
@@
Figure 3: Left: SuperCube illuminated with UV light prior to completion. The absence of a white reflector layer on the SuperCube’s top surface provides a clear view of the inner structure, featuring optically isolated PS cubes intersected by WLS fibers. Right: FIM manufactured 5x5x5 SuperCube, requiring no post-processing.
@@
The response of the SuperCube was characterized with cosmic particles in terms of single-cube scintillation light yield and cube-to-cube scintillation light crosstalk, which are the main parameters that determine the detector performance in terms of particle tracking, identification and calorimetry. Two detected cosmic particle events are shown in Fig. 4. Particles producing a vertical track, whose range in the detector can be more precisely estimated, were selected and used for analysis. Being mostly minimum ionizing particles (MIP), their typical energy deposited in PS is approximately 1.8 MeV/cm, which allows to precisely estimate the number of photons detected per unit energy loss. Moreover, with an energy spectrum that spans from a few hundred MeV to several GeV, cosmic particles typically produce through-going tracks in the detector with a very uniform energy deposition. By measuring the number of PE across all the readout channels and geometrically matching the two detector projections of the event, the 3D track of the particle can be reconstructed and the number of scintillation photons produced in each cube can be accurately measured. As shown in Fig. 5, the scintillation light yield of the 3D printed prototype was found to be comparable to that obtained with the one produced by cast polymerization (using the same scintillator composition) in [41]. A most probable value (MPV) of about 29 PE was measured for the 3D printed SuperCube, close to the MPV of the prototype produced with cast polymerization (about 28 PE).
Figure 4: Top: FIM manufactured SuperCube prototype instrumented with WLS fibers and SiPMs. Bottom: events of two detected cosmic particles crossing the SuperCube from top to bottom with a detected diagonal (left) and vertical (right) track. The 2D projections show the number of photons detected in each readout channel (color scale), while the 3D voxels show the reconstructed 3D particle track. In the matching of the 2D projections channels with less then 7 PE, likely from cube-to-cube light crosstalk, were not considered. The red line indicates the fitted particle trajectory.
@@ III DISCUSSION
The FDM manufactured PS showed a sufficient transparency for centimeter-sized voxels, but the FIM fabricated samples visibly improved the quality by reducing imperfections created by the merger of different deposited lines during the FDM process and the reduction of air bubbles within the PS volume. The improved transparency results from the straight down-to-up filling pattern, which creates a single melt pool that expands and forms the PS structure within the optical reflective frame. The nozzle’s orifice remained submerged in the melt pool until the PS forming process was complete. This technique prevented air from being trapped between different layers of melted polymer. Additionally, the forming of a cubic-centimeter sized PS cube with FIM needs around 30 seconds, whereas the same structure takes around 5 minutes using FDM fabrication, i.e., about ten times as long. The faster per-unit manufacturing time is crucial when it comes to the production of large volume, highly-segmented particle detectors that can include millions of isolated voxels.
@@
The geometrical tolerance of the optically isolating frame produced by FIM is improved compared to the particle detector manufactured solely by FDM. This is a result of the introduction of a heat-resistant material that is able to preserve its designed shape during the filling of melted PS, thus creates a distinct boundary between the individual PS voxels. In FDM, both the scintillation and the reflective filament are made of plastics with similar thermal characteristics, which can result in a mixing at the boundary layers between these two materials and also a warping and distortion of the reflective geometry. Although the tolerance cannot be considered as good as the one achievable with injection molding or cast polymerization (better than $50~{}\mu\text{m}$ for plastics), it is worth noting that it is affected by the type of 3D printed material, the printing strategy and the precision of the machine, which will be improved in future studies.
@@
The largest advantage of the FIM method compared to FDM or more traditional methods is the capability of creating precise, close-fitting holes through the entire particle detector to host WLS fibers. FDM is notorious for producing poor quality small holes in terms of size consistency, shape and surface roughness, which results in a low WLS fiber to PS contact area, especially perpendicularly to the print plane. Injection molding has the capability to create hollow components within the manufactured structure. However, incorporating a white reflector without causing any harm to the pre-existing holes poses a significant challenge. In our next prototypes also a third vertical hole will be added. An alternative option would be to utilize thin pipes capable of withstanding temperatures significantly higher than 200°C, in place of metal rods. This modification would facilitate the easy insertion of WLS fibers, leading to the creation of the SuperCube. A single-layer prototype was successfully produced using borosilicate glass pipes with outer and inner diameters of 1.5 mm and 1 mm, respectively. While this approach eliminates the need to extract rods after the injection and solidification of the plastic scintillator, the addition of non-plastic material does slightly increase the inactive volume. Consequently, we opted to prioritize the baseline option involving metal rods.
@@
In terms of the detection performance, the quantity of scintillation light, collected from the FIM manufactured PS cubes crossed by cosmic particles, is similar to that observed from PS cubes produced with traditional methods like cast polymerization. The cube-to-cube crosstalk at the few-percent level is comparable to that of typical PS detectors [30] and is low enough to provide unambiguous 3D particle tracking.
@@
The light crosstalk is about four times higher than that with the cast polymerization prototype, due to the higher wall transmittance. Hence, one would expect an approximately 20% lower total light yield. However, this was not observed, as shown in Fig. 5. One may conclude that the additive manufactured PS has an intrinsic light output higher than the one from cast polymerization, but this was not the result of our past work [55, 53]. It is well known that the light yield of a PS detector can be enhanced by improving the light trapping efficiency of the WLS fibers through the reduction of air gaps, that is the increase of the contact area between the PS and the WLS fibers, whose refractive indices are very similar and thus allow a high transmittance from the scintillation material to the fiber. This can be obtained by using optical grease [59, 60, 61], which is difficult to homogeneously introduce in long, small-diameter holes. Using the FIM method, it was possible to create holes, that span the length of the SuperCube, of only 100 $\mu$m larger than the WLS fibers, resulting in a large contact surface between the PS cubes and the fibers. This feature is expected to increase the total light yield and to compensate for the light loss from the observed crosstalk.
A preliminary time resolution measurement was conducted using a Hamamatsu H6410 photomultiplier tube (PMT) coupled to a sample of 3D printed plastic scintillator exposed to a ${}^{60}\text{Co}$ radioactive source. The data was read out with a LeCroy Wave runner 104MXi-A oscilloscope, revealing a decay time consistent with that of UPS-923A, indicating no degradation of the timing properties resulting from the scintillator additive manufacturing process. However, a more comprehensive quantitative characterization is deferred to future work and publication.
Aging presents another critical concern, potentially resulting in a few percent reduction in light yield annually for extruded polystyrene-based scintillators [56, 62, 63, 64]. Given the similarity of the FIM process to extrusion or injection molding in the treatment of the plastic scintillator material, a similar degradation is expected. Despite various measurements conducted over approximately four months, no visible changes in light yield and crosstalk were observed within the precision of the experimental setup. Nevertheless, more precise quantitative results require dedicated aging measurements, which are planned for future investigation.
@@
In conclusion, we have successfully achieved the additive manufacturing of a 3D optically-segmented plastic scintillator detector. To the best of our knowledge, this is the first time such a feat has been accomplished. This detector is capable of both particle tracking and energy loss measurement. It demonstrates a tool that opens new possible concepts of particle detectors characterized by a very fine 3D granularity in very large active volumes, above the tonne-scale. Although not tested in this work, millimeter-sized granularity and bigger units, e.g. $30\times 30\times 30~{}\text{cm}^{3}$ or above, are expected to be achieved without problems. We believe that this new technology opens up new possibilities for the future of particle physics experiments. In particular, long-baseline accelerator neutrino oscillations [65, 66, 67, 68] or those searching for new neutrino sterile states via short-baseline oscillations [69, 70, 71, 72, 73, 74] can profit of the developed additive manufacturing process by building very large but finely-segmented scintillator detectors to obtain a large sample of very detailed images of neutrino interactions. Moreover, the developed production method will allow to easily build highly-performing fine-granularity electromagnetic or hadronic sampling calorimeters, enabling high-resolution particle flow analysis and fulfill the requirements of future collider experiments [31, 75]. FIM can be performed with various types of filaments, where the white reflector could be covered by an additional layer made of heavier nuclei or doped material. Finally, large-volume 3D detector geometries, made possible by the FIM process, can allow for a high-efficiency detection of fast neutrons with a precise measurement of their time of flight, thanks to the large content of hydrogen in PS and its short decay time [28, 35]. This is of fundamental importance for next generation neutrino physics experiments, providing constraints on leading systematic uncertainties [34, 76, 33, 77].